OpenStudy (anonymous):

There are 20 terms in an A.P. the sum of the first 10 terms is 55 and the sum of the last 10 terms is 355. Find first term and common difference.

OpenStudy (anonymous):

S10 = 55 = (10/2)[2a1 + 9d] S20 = 410 = (20/2)[a1 + 19d] (S20 = 55 + 355) 5(2a1 + 9d) = 55 10(2a1 + 19d] = 410 solving these 2 equations will give you first term a1 and common difference d

OpenStudy (anonymous):

can u solve these?

OpenStudy (anonymous):

Hey I tried to solve it but the answer is not correct!

OpenStudy (anonymous):

Oops, my bad! the answer is correct. Sorry and thanks alot! :)